# RGS-DR: Reflective Gaussian Surfels with Deferred Rendering for Shiny Objects

### v3

@@ Abstract
We introduce RGS-DR, a novel inverse rendering method for reconstructing and rendering glossy and reflective objects with support for flexible relighting and scene editing. Unlike existing methods (e.g., NeRF and 3D Gaussian Splatting), which struggle with view-dependent effects, RGS-DR utilizes a 2D Gaussian surfel representation to accurately estimate geometry and surface normals, an essential property for high-quality inverse rendering. Our approach explicitly models geometric and material properties through learnable primitives rasterized into a deferred shading pipeline, effectively reducing rendering artifacts and preserving sharp reflections. By employing a multi-level cube mipmap, RGS-DR accurately approximates environment lighting integrals, facilitating high-quality reconstruction and relighting. A residual pass with spherical-mipmap-based directional encoding further refines the appearance modeling. Experiments demonstrate that RGS-DR achieves high-quality reconstruction and rendering quality for shiny objects, often outperforming reconstruction-exclusive state-of-the-art methods incapable of relighting.
@@ 1 Introduction
To address these issues, inverse neural rendering aims to decompose scene properties and approximate light transport. Techniques in this area [33, 12, 16] have led to significant improvements in rendering shiny objects. A core component of these methods is the estimation of reflection directions based on surface normals, which is essential for modeling specular reflections under environmental lighting. The quality of reconstruction and rendering is highly dependent on the accuracy of these surface normals. To obtain them, Ref-NeRF [33] uses multilayer perceptrons to learn coordinate-based normals, while GaussianShader [12] employs learnable normal residuals conditioned on the viewing direction. However, both methods detach the normal estimation from the underlying geometry, which may result in noisy normals. Other methods, such as NMF [21] and ENVIDR [17], achieve better performance by using normals derived from the gradient of the geometry, while 3DGS-DR [16] relies on the shortest axis of each Gaussian ellipsoid to estimate normals that may or may not align with the actual surface of the object. In contrast, Ref-GS [39] takes advantage of standard surfel modeling [25], adopting 2D-oriented disks as surface elements, providing an explicit normal representation that enhances the modeling of shiny objects. However, Ref-GS encodes some of the object’s material properties implicitly together with the ambient lighting, making it difficult to edit the scene, such as changing the environment lighting.
@@
To achieve relighting with inverse rendering, it is essential to explicitly model the environment lighting within the rendering equation. This explicit modeling enables the seamless replacement of the original lighting with the desired target lighting for effective relighting. Existing methods approximate the rendering equation using different formulations to enable this explicit modeling. A key challenge in this process is the accurate modeling of the integral of incident light from different directions. Spherical Harmonics [21] and Spherical Gaussians [33, 37] are commonly used to represent the distribution of environment lighting, as they provide closed-form solutions for approximating light integration. Other methods [12, 16, 39] utilize discrete grid textures to capture high-frequency signals in environment lighting. However, 3DGS-DR [16], the state-of-the-art method in this domain, fails to properly formulate the integral using only a single-level environment map, which limits its expressive capability. Meanwhile, GaussianShader [12], which samples from multi-level mipmaps to approximate the integral, struggles with geometric uncertainty, resulting in degraded results.
In this paper, we introduce RGS-DR, a method that enables high quality shiny object reconstruction and rendering, while supporting scene relighting. RGS-DR inherits the surfel representation from 2DGS [11], enabling accurate geometry modeling and surface normal computation. We decompose the scene’s geometric and material properties and store them as learnable variables within each 2D Gaussian primitive. In the proposed rendering pipeline, the properties of these Gaussians are rasterized to generate a material property buffer in image space. We use a deferred rendering method, applying shading functions to each pixel in the graphics buffer to determine its final color. This approach mitigates the impact of depth precision on rendering and prevents the degradation of rendering quality due to the accumulation of geometric errors and uncertainties during alpha blending - an issue commonly found in traditional methods. In these shading functions, the light transport is considered, and the environment map is explicitly modeled using mip-mapping. We use a multi-level cubemap structure that enhances the accuracy of the light integral approximation on a discrete grid. We also model inter-reflections as a residual in a deferred rendering formulation and incorporate a residual pass that includes the spherical encoding from Ref-GS [39] that further enhances the rendering quality. Additionally, our method produces relighting effects by replacing the content of the explicit environment map, while modifying the Gaussians’ material attributes—such as roughness and diffuse color—unlocks scene editing capabilities. Experimental results show that RGS-DR achieves high reconstruction quality, on par with state-of-the-art reconstruction-only techniques, while also providing plausible estimates of the environment illumination and the material properties of the scene.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [15] has recently emerged as a promising alternative to NeRF for 3D scene representation, delivering high-quality novel view synthesis and rapid rendering speeds. Advancements in 3DGS, such as the use of structured Gaussians [20, 6, 26], texture mapping [5], gradient cues [40, 27], and anti-aliasing [36], have further enhanced rendering quality. However, these methods often fall short in accurately reconstructing and rendering glossy or reflective surfaces due to their inability to precisely model surface-light interactions. To address these limitations, 2D Gaussian Splatting (2DGS) [11] projects Gaussian disks onto object surfaces and applies local smoothing. This approach ensures view-consistent geometry. Our method leverages its explicit surface normals combined with a tailored shading function to accurately model light transport in scenes featuring shiny objects.
Inverse Neural Rendering. Inverse rendering is a challenging problem that seeks to reconstruct scene properties, including geometry, materials, and lighting, from images. Recent advances in inverse rendering have integrated differentiable volume rendering with implicit neural representations, drawing inspiration from NeRF [22], to learn scene representations encompassing both geometry and materials. Early approaches in this area often made strong assumptions or required additional priors, such as a known illumination [29, 2], geometry [38], or varying illumination conditions [2, 10]. These limitations have been addressed in subsequent works [37, 3, 4, 17, 9, 13, 33], which jointly estimate geometry, materials (e.g., BRDF), and lighting. Gaussian-based neural rendering methods [12, 32, 35, 41] provide faster solutions. These studies calculate reflection directions to model light transport, thereby achieving superior performance compared to traditional neural rendering methods when handling shiny objects. Recent works [39, 16] on this task employ deferred rendering [8] on material buffer images. Our method inherits the deferred rendering scheme and explicitly models both the material properties and environment map without relying on alpha blending. This approach produces sharp reflection effects and facilitates property editing, as illustrated in Fig. 1.
@@ 4 Methodology
Figure 2: Our rendering pipeline consists of three passes. The geometry pass produces screen-space diffuse color $\mathcal{I}_{d}$, specular tint $S$, roughness $M$, normals $N$, and low-dimensional features $K$, which feed into the subsequent passes. The lighting pass employs a cube mipmap to model environmental light for shading as described in [12]. Meanwhile, the residual pass uses a spherical-mip-based directional encoding (inspired by [39]) along with a shallow MLP $f_{r}$ to predict view-dependent effects not captured by the lighting pass.
@@
Overview. Instead of first calculating the view-dependent color of primitives and then blending them, we employ a deferred rendering scheme. Our approach rasterizes the geometry and blends the properties into a Graphics Buffer (G-Buffer), which stores buffer images for each property (§ 4.1). The G-Buffer stores essential information for modeling light transport in reflections, including geometry, material, and lighting-related properties. It enables a representation of global illumination effects. Leveraging this data, shading functions are applied to approximate the rendering equation based on the material attributes in the G-Buffer. This process enables the computation of pixel-wise colors for the specular component while accounting for indirect lighting contributions from surrounding environment (§ 4.2). In addition to environmental illumination, light undergoes multiple reflections between object surfaces, a phenomenon known as inter-reflection, which is not explicitly modeled in standard shading functions. To address this limitation, we propose a residual rendering pass that integrates features from the G-Buffer with a spherical feature mip-map  to effectively encode inter-reflection effects (§ 4.3). We integrate these components into our proposed differentiable pipeline to generate the final rendered images. The complete pipeline is illustrated in Fig. 2.
Deferred Rendering. RGS-DR approximates the rendering equation with:
$\mathbf{c}(\boldsymbol{\omega}_{o})=\mathbf{c}_{d}+\mathbf{s}\odot\mathbf{L}_{s}(\boldsymbol{\omega}_{o},\mathbf{n},\rho)+\mathbf{c}_{r}(\boldsymbol{\omega}_{o}),$ (3)
@@ 4.1 G-Buffer Rendering
We adapt the alpha blending formulation in Eq. 1 to generate buffer images by replacing $\mathbf{c}_{i}$ with the corresponding property vectors. RGS-DR fills the G-buffer with the diffuse map $\mathbf{I}_{d}$, roughness map $\mathbf{M}$, specular tint map $\mathbf{S}$, feature map $\mathbf{K}$, and normal map $\mathbf{N}$, as well as the accumulated alpha map $\mathbf{A}$. They have the same resolution as the render target’s but differ in the number of channels. Most of them are alpha-blended from corresponding properties stored in each Gaussian primitive except for the normal map $\mathbf{N}$, which is derived from the primitive geometry using Eq. 2. Given a specific pixel coordinate, the desired property at that location can be obtained by interpolating the corresponding values from the buffer image.
@@ 4.2 Deferred Shading Functions
Figure 3: Specular light computation for the lighting pass using a differentiable environment light encoded as a cube mipmap.
@@
[…]
where $\Omega$ denotes the directions of the entire upper hemisphere above the current primitive disk, and $D$ is the GGX [34] normal distribution function; $\boldsymbol{\omega}_{i}$ is the direction of the input radiance.
The integration is approximated using the split-sum approximation to avoid computationally expensive explicit integral evaluations. Specifically, we achieve this using a pre-filtered cube mip-map $\mathbf{E}$ where each mip-map level stores progressively blurrier versions of the original environment map, simulating different levels of surface roughness for reflections. Paired with a Look-Up Table $\digamma$ for Fresnel and Geometric terms [7] that modulates specular reflections, we approximate the integral calculation using following formulation:
$\mathbf{L}_{s}(\boldsymbol{\omega}_{o},\mathbf{n},\rho)=~{}$ $\mathbf{s}\cdot\digamma(\mathbf{s},\rho,\mathbf{n},\boldsymbol{\omega}_{o})\cdot\Phi(\boldsymbol{\omega}_{r},\rho,\mathbf{E})$
$+(1-\mathbf{s})\cdot\Phi(\boldsymbol{\omega}_{r},\rho_{0},\mathbf{E}),$ (6)
where the reflection direction is given by $\boldsymbol{\omega}_{r}=2\bigl{(}\boldsymbol{\omega}_{o}\cdot\mathbf{n}\bigr{)}\mathbf{n}-\boldsymbol{\omega}_{o}$. The function $\Phi(\cdot)$ samples the mip-map $\mathbf{E}$ at a specific level determined by the input roughness. Here, $\rho_{0}$ represents a high roughness value. Consequently, $\Phi(\boldsymbol{\omega}_{r},\rho_{0},\mathbf{E})$ characterizes the ambient lighting from the environment. It enhances the low-frequency reflection. Figure 3 illustrates the pipeline of computing $\mathbf{L}_{s}$.
@@ 4.3 Residual Rendering Pass
For a given screen-space point $\mathbf{u}(u,v)$, the spherical mip encoding can be calculated based on the viewing direction $\boldsymbol{\omega}_{o}$, normal $\mathbf{n}$, roughness $\rho$, and feature $\mathbf{k}$. The viewing direction and normal are used to estimate the reflected direction $\boldsymbol{\omega}_{r}$ which is then converted into polar coordinate $(\theta_{r},\phi_{r})$. The spherical mip encoding is then computed by trilinear interpolation on the spherical mipmap $\mathcal{M}$ along the roughness dimension and given by $\mathbf{h}=\mathrm{Sph\text{-}Mip}\bigl{(}\boldsymbol{\omega}_{r},\rho,\mathcal{M}\bigr{)}$. This feature is concatenated with the outer product $\mathbf{h}\otimes\mathbf{k}$ to form the directional encoding, which is used as input to a shallow MLP $f_{r}$ to produce the residual color term $\mathbf{c}_{r}=f_{r}(\mathbf{h},\mathbf{k}\otimes\mathbf{h})$ and the corresponding residual image
$\mathcal{I}_{r}=\textit{f}_{r}(\mathbf{H},\mathbf{K}\otimes\mathbf{H}).$ (7)
@@ 4.4 Training Strategy
Furthermore, we use an alpha loss to prevent floaters caused by redundant Gaussians:
$\mathcal{L}_{\mathrm{\alpha}}={\bigl{\lVert}\hat{A}-A\bigr{\rVert}}_{1},$ (11)
where $\hat{A}$ is the accumulated alpha map in the G-buffer and $A$ is the corresponding ground-truth alpha channel of the provided RGBA image.
@@
where $\lambda_{d}$ is the weight of the depth distortion loss, $\lambda_{n}$ is the weight of the normal consistency loss and $\lambda_{\alpha}$ is the weight of the silhouette loss.

### v4

# RGS-DR: Deferred Reflections and Residual Shading in 2D Gaussian Splatting

@@ Abstract
In this work, we address specular appearance in inverse rendering using 2D Gaussian splatting with deferred shading and argue for a refinement stage to improve specular detail, thereby bridging the gap with reconstruction-only methods. Our pipeline estimates editable material properties and environment illumination while employing a directional residual pass that captures leftover view-dependent effects for further refining novel view synthesis. In contrast to per-Gaussian shading with shortest-axis normals and normal residuals, which tends to result in more noisy geometry and specular appearance, a pixel-deferred surfel formulation with specular residuals yields sharper highlights, cleaner materials, and improved editability. We evaluate our approach on rendering and reconstruction quality on three popular datasets featuring glossy objects, and also demonstrate high-quality relighting and material editing.
@@ 1 Introduction
To address these issues, inverse neural rendering aims to decompose scene properties and approximate light transport. Techniques in this area [35, 22, 18, 12, 16, 43, 38] have led to significant improvements in rendering shiny objects. A core component of these methods is the estimation of reflection directions based on surface normals, which is essential for modeling specular reflections under environmental lighting. The quality of reconstruction and rendering is highly dependent on the accuracy of these surface normals. To obtain them, Ref-NeRF [35] uses multilayer perceptrons to learn coordinate-based normals, while GaussianShader [12] employs learnable normal residuals conditioned on the viewing direction. However, both methods detach the normal estimation from the underlying geometry, which may result in noisy normals. Other methods, such as NMF [22] and ENVIDR [18], achieve better performance by using normals derived from the gradient of the geometry, while 3DGS-DR [16] relies on the shortest axis of each Gaussian ellipsoid to estimate normals that may or may not align with the actual surface of the object. In contrast, Ref-GS [43] takes advantage of standard surfel modeling [26], adopting 2D-oriented disks as surface elements, providing an explicit normal representation that enhances the modeling of shiny objects. However, Ref-GS [43] encodes some of the object’s material properties implicitly together with the ambient lighting, thus preventing relighting or other scene editing tasks.
@@
To achieve relighting with inverse rendering, it is essential to explicitly model the environment lighting within the rendering equation. This explicit modeling enables the seamless replacement of the original lighting with the desired target lighting for effective relighting. Existing methods approximate the rendering equation using different formulations to enable this explicit modeling. A key challenge in this process is the accurate modeling of the integral of incident light from different directions. Spherical Harmonics [22] and Spherical Gaussians [35, 41] are commonly used to represent the distribution of environment lighting, as they provide closed-form solutions for approximating light integration. Other methods [12, 16, 38] utilize discrete grid textures to capture high-frequency signals in environment lighting. However, 3DGS-DR [16] fails to properly formulate the integral using only a single-level environment map and does not learn explicit material properties, which limit its expressive capability and editability. Meanwhile, GaussianShader [12], which samples from multi-level mipmaps to approximate the integral, struggles with geometric uncertainty, resulting in degraded results. Ref-Gaussian [38] reduces geometric uncertainty by utilizing the shortest axes of 2D Gaussians, but struggles to decompose diffuse from specular appearance and has to fit the diffuse color with spherical harmonics, disentangled from base color, leading to incorrect estimation of the latter.
In this paper we present RGS-DR, an inverse-rendering method for glossy objects explicitly scoped to direct image-based lighting (IBL). As in 2DGS [11] and related work [43, 38, 34], we use surfels to obtain reliable geometry and normals. Scene geometry and physically based rendering (PBR) parameters are optimized on each 2D Gaussian and rasterized into an image-space material buffer via a pixel-deferred pipeline. Per-pixel shading with split-sum IBL improves the discrete approximation of the specular integral and avoids per-Gaussian alpha-blend artifacts typical of forward shading [12]. To better capture view-dependent detail, we add a lightweight directional residual in image space (Ref-GS encoding [43]) that fits leftover specular effects such as glints and micro-geometry. Relighting and scene edits are then performed by swapping the estimated environment map and adjusting the recovered material maps (e.g., roughness, diffuse albedo). Empirically, RGS-DR achieves competitive reconstruction under training illumination while providing consistent environment and material estimates, and it supports high-quality relighting and editing on datasets with highly reflective objects. Our contributions can be summarized as follows:
  • A 2DGS-based, inverse-rendering pipeline with deferred reflections that recovers editable materials and environment illumination.
  • A refinement stage with a directional residual module that sharpens specular appearance for novel view synthesis and reduces the gap to reconstruction-only baselines.
  • A detailed evaluation of rendering, reconstruction, relighting, and material editing, compared to (direct IBL) inverse rendering and reconstruction-only baselines.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [15] has recently emerged as a promising alternative to NeRF for 3D scene representation, delivering high-quality novel view synthesis and rapid rendering speeds. Advancements in 3DGS, such as the use of structured Gaussians [21, 6, 27], texture mapping [5], gradient cues [44, 28], and anti-aliasing [40], have further enhanced rendering quality. However, these methods often fall short in accurately reconstructing and rendering glossy or reflective surfaces due to their inability to precisely model surface-light interactions. To address these limitations, 2D Gaussian Splatting (2DGS) [11] projects Gaussian disks onto object surfaces and applies local smoothing. This approach ensures view-consistent geometry. Our method and other recent methods [43, 38, 34] leverage its explicit surface normals combined with a tailored shading function to accurately model light transport in scenes featuring shiny objects.
Inverse Neural Rendering. Inverse rendering is a challenging problem that seeks to reconstruct scene properties, including geometry, materials, and lighting, from images. Recent advances in inverse rendering have integrated differentiable volume rendering with implicit neural representations, drawing inspiration from NeRF [23], to learn scene representations encompassing both geometry and materials. Early approaches in this area often made strong assumptions or required additional priors, such as a known illumination [30, 2], geometry [42], or varying illumination conditions [2, 10]. These limitations have been addressed in subsequent works [41, 3, 4, 18, 9, 13, 35, 20], which jointly estimate geometry, materials (e.g., BRDF), and lighting.
Gaussian-based neural rendering methods [12, 33, 36, 45] have emerged as faster, more specular-aware alternatives to field-based NeRFs by explicitly computing reflection directions and using BRDF-aware shading. Several recent works [43, 16, 38, 34] adopt a deferred formulation [8], rasterizing material buffers and shading in image space. We follow this pixel-deferred paradigm and explicitly estimate PBR material maps together with an HDR environment, performing split-sum IBL per pixel. This avoids per-primitive alpha-blended shading, which tends to smear speculars and couple errors to depth ordering (see Fig. 1). Closest to our setting, GaussianShader [12] attaches a per-Gaussian residual color modeled with low-order spherical harmonics to absorb unmodeled view dependence. In contrast, we employ a deferred residual pass with the more expressive spherical directional encoding of Ref-GS [43]. Residuals are predicted in image space (not stored per primitive), so we keep only a small per-Gaussian feature vector (much smaller than SH) for conditioning, yielding fewer parameters and lower memory at comparable cost. The residual is disabled at relight time to prevent training-illumination bias. Several methods [20, 38, 39, 7] extend split-sum IBL with additional components for indirect lighting, aiming to capture complex global illumination effects. However, these enhancements fall outside the scope of direct IBL, which is our focus. Other works [17, 34, 37] adopt fundamentally different assumptions, such as multi-light supervision or external geometric priors, making them incompatible with our setting. These directions target complementary problems and are not directly comparable to ours.
@@ 4 Methodology
Figure 2: Our rendering pipeline consists of three passes. The geometry pass produces screen-space diffuse color $\mathcal{I}_{d}$, specular tint $S$, roughness $M$, normals NN, and low-dimensional features $K$, which feed into the subsequent passes. The lighting pass employs a cube mipmap to model environmental light for shading as described in [12]. Meanwhile, the residual pass uses a spherical-mip-based directional encoding (inspired by [43]) along with a shallow MLP $f_{\text{res}}$ to predict view-dependent effects not captured by the lighting pass.
@@
Overview. Instead of first calculating the view-dependent color of primitives and then blending them, we employ a deferred rendering scheme. Our approach rasterizes the geometry and blends the properties into a Graphics Buffer (G-Buffer), which stores buffer images for each property (§ 4.1). The G-Buffer stores essential information for modeling light transport in reflections, including geometry, material, and lighting-related properties. It enables a representation of global illumination effects. Leveraging this data, shading functions are applied to approximate the rendering equation based on the material attributes in the G-Buffer. This process enables the computation of pixel-wise colors for the specular component while accounting for indirect lighting contributions from the surrounding environment (§ 4.2). In addition to environmental illumination, light undergoes multiple reflections between object surfaces, a phenomenon known as inter-reflection, which is not explicitly modeled in standard shading functions. To address this limitation, we propose a residual rendering pass that integrates features from the G-Buffer with a spherical feature mip-map [43] to effectively encode inter-reflection effects (§ 4.3). We integrate these components into our proposed differentiable pipeline to generate the final rendered images. The complete pipeline is illustrated in Fig. 2.
Deferred Rendering. RGS-DR adopts the rendering equation approximation of GaussianShader [12] defined as:
$\mathbf{c}(\boldsymbol{\omega}_{o})=\mathbf{c}_{d}+\mathbf{s}\odot\mathbf{L}_{s}(\boldsymbol{\omega}_{o},\mathbf{n})+\mathbf{c}_{r}(\boldsymbol{\omega}_{o}),$ (3)
@@ 4.1 G-Buffer Rendering
We adapt the alpha blending formulation in Eq. 1 to generate buffer images by replacing $\mathbf{c}_{i}$ with the corresponding property vectors. Following the notation of Ref-GS[43], RGS-DR fills the G-buffer with the diffuse map $\mathcal{I}_{d}$, roughness map $\mathbf{M}$, specular tint map $\mathbf{S}$, feature map $\mathbf{K}$, and normal map $\mathbf{N}$, as well as the accumulated opacity $\mathbf{A}$. They have the same resolution as the render target but differ in the number of channels. Most of them are alpha-blended from corresponding properties stored in each Gaussian primitive except for the normal map $\mathbf{N}$, which is derived from the primitive geometry using Eq. 2.
@@ 4.2 Deferred Shading Functions
$\mathbf{L}_{s}(\boldsymbol{\omega}_{o},\mathbf{n})=\int_{\Omega}L(\boldsymbol{\omega}_{i})\,f_{r}(\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o})\,\bigl(\boldsymbol{\omega}_{i}\cdot\mathbf{n}\bigr)\,d\boldsymbol{\omega}_{i},$ (5)
where $\boldsymbol{\omega}_{i}$ is the direction of the input radiance, $\Omega$ denotes the directions of the entire upper hemisphere above the current primitive disk, and $f_{r}$ is the bidirectional reflectance distribution function.
We use the split-sum approximation of the integral to avoid computationally expensive explicit integral evaluations. Specifically, we achieve this using a pre-filtered cube mip-map $\mathbf{E}$ where each mip-map level stores progressively blurrier versions of the original environment map, simulating different levels of surface roughness for reflections.
@@ 4.3 Residual Rendering Pass
For a given screen-space point $\mathbf{u}(u,v)$, the spherical mip encoding can be calculated based on the viewing direction $\boldsymbol{\omega}_{o}$, normal $\mathbf{n}$, roughness $\rho$, and feature $\mathbf{k}$. The viewing direction and normal are used to estimate the reflected direction $\boldsymbol{\omega}_{r}$ which is then converted into polar coordinate $(\theta_{r},\phi_{r})$. The spherical mip encoding is then computed by trilinear interpolation on the spherical mipmap $\mathcal{M}$ along the roughness dimension and given by $\mathbf{h}=\mathrm{Sph\text{-}Mip}\bigl(\boldsymbol{\omega}_{r},\rho,\mathcal{M}\bigr)$. This feature is concatenated with the outer product $\mathbf{h}\otimes\mathbf{k}$ to form the directional encoding, which is used as input to a shallow MLP $f_{\text{res}}$ to produce the residual color term $\mathbf{c}_{r}=f_{\text{res}}(\mathbf{h},\mathbf{k}\otimes\mathbf{h})$ and the corresponding residual image
$\mathcal{I}_{r}=\textit{f}_{\text{res}}(\mathbf{H},\mathbf{K}\otimes\mathbf{H}).$ (6)
@@ 4.4 Training Strategy
Furthermore, we use an alpha loss to prevent floaters caused by redundant Gaussians in synthetic scenes:
$\mathcal{L}_{\mathrm{\alpha}}={\bigl\lVert A-A_{gt}\bigr\rVert}_{1},$ (10)
where $A$ is the accumulated alpha map in the G-buffer and $A_{gt}$ is the corresponding ground-truth alpha channel of the provided RGBA image.
@@
where $\lambda_{d}$, $\lambda_{n}$, and $\lambda_{\alpha}$ are the corresponding loss weights.